OpenStudy (anonymous):

Elaine is painting a fence. She knows that she needs 1 gallon of paint for each 10 feet of fencing. The fencing will enclose a rectangular yard that is 3 yards long and 2 yards wide. How much paint does Elaine need? (Remember 3 feet = 1 yard)

OpenStudy (anonymous):

First find the perimeter. \[P = 2(l + w) = ?\]

OpenStudy (anonymous):

i ended up with 6

OpenStudy (anonymous):

2(2 + 3) = 2(5) = ?

OpenStudy (anonymous):

i ended up with 2. i think im calculating all wrong

OpenStudy (anonymous):

2(5) is the same as 2 x 5.

OpenStudy (anonymous):

its 10

OpenStudy (anonymous):

Ok. Then multiply 10 by 3 to convert it from yards to feet.

OpenStudy (anonymous):

i got 30

OpenStudy (anonymous):

she would need 3 gallons right

OpenStudy (anonymous):

Um sorrry...I got distracted. Yes. You are right :)